OpenStudy (anonymous):

help plsssss... medals and fan solve dx/dt = 0.05 (20-x)

OpenStudy (anonymous):

@sammixboo help pls ?

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@IrishBoy123

OpenStudy (anonymous):

@dan815

OpenStudy (irishboy123):

@yajna how did you try it? useful to know how experienced you are at this.

OpenStudy (anonymous):

by integration, differential

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

\[\Large \frac{dx}{dt} = 0.5(20-x)\] \[\Large dx = 0.5(20-x)*dt\] \[\Large \frac{dx}{20-x} = 0.5dt\] \[\Large \int\frac{dx}{20-x} = \int 0.5dt\] I'll let you finish up

OpenStudy (anonymous):

yeah i ve reached so far,, i am having issues

jimthompson5910 (jim_thompson5910):

hint: let u = 20 - x, so du = -dx which means dx = -du

jimthompson5910 (jim_thompson5910):

\[\Large \int\frac{dx}{20-x} = \int 0.5dt\] \[\Large \int\frac{1}{20-x}dx = \int 0.5dt\] \[\Large \int\frac{1}{u}(-du) = \int 0.5dt\] \[\Large -\int\frac{1}{u}du = \int 0.5dt\]

OpenStudy (anonymous):

so far i have done it right

jimthompson5910 (jim_thompson5910):

ok what's next?

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (freckles):

what part you need help on?

OpenStudy (freckles):

and up to what do you have?

OpenStudy (anonymous):

I HAVE TO INTEGRATE IT

OpenStudy (freckles):

yep can you do that?

OpenStudy (freckles):

these integrals you have are pretty easy to evaluate (no sub needed ) \[\int\limits_{}^{}\frac{1}{x} dx=\ln|x|+C \\ \int\limits_{}^{}c dx=cx+K\]

OpenStudy (anonymous):

WHAT about 1/(20-x) ?

OpenStudy (freckles):

@jim_thompson5910 took the liberty of doing the sub for you

OpenStudy (freckles):

you do know that if u=20-x then du/dx=-1 or du=- dx right?

OpenStudy (freckles):

 that is how he got the integral on the left hand side